OpenStudy (anonymous):

2√(32x^2)-8x√(8)-3√(18x^2) combine the expressions

OpenStudy (saifoo.khan):

C'mon Myininaya!

myininaya (myininaya):

\[2\sqrt{32x^2}-8x \sqrt{8}-3\sqrt{18x^2}=2\sqrt{x^2}\sqrt{16}\sqrt{2}-8x \sqrt{4}\sqrt{2}-3\sqrt{x^2}\sqrt{9}\sqrt{2}\] \[=2*x*4*\sqrt{2}-8*x*2*\sqrt{2}-3*x*3*\sqrt{2}=\sqrt{2}(8x-16x-9x)=-17x \sqrt{2}\]

myininaya (myininaya):

oops it got cut off

myininaya (myininaya):

its suppose to say at the end \[-17x \sqrt{2}\]